This one is for d genius.
A girl went to school & the teacher asked how old are you? She replied I don't know, bt I know that I'm half my mother's age. The teacher asked ok how old is your mother? She said I don't know but I know that my Dad is 5 yrs older than my mother. Teacher, how old is yor dad? She said I don't know but I know that if we add our ages the answer is 100 yrs... Nw how old is the girl, her mum her dad?

Answer by ankor@dixie-net.com(22740)

Just get rid of the nonsense, and write an equation for each phrase
:
I'm half my mother's age.
g = .5m
my Dad is 5 yrs older than my mother.
f = m + 5
If we add our ages the answer is 100 yrs...
f + m + g = 100
replace g with .5m
replace f with (m+5)
(m+5) + m + .5m = 100
2.5m = 100 - 5
2.5m = 95
m = 95/2.5
m = 38 yrs is Mom's age
then
g = .5(38)
g = 19 is girl's age
and
f = 38 + 5
f = 43 is Dad's age
:
:
See if that adds up
19 + 38 + 43 = 100